# Energy solutions of the Cauchy–Dirichlet problem for fractional nonlinear diffusion equations

### v1

@@ 1. Introduction
Most of studies on fractional nonlinear diffusion equations have been performed for the Cauchy problem, where the equations are posed in $\mathbb{R}^{d}$ (see, e.g., [34, 35, 64, 20, 32, 33, 14] and references therein). On the other hand, there have also been a limited number of contributions to the Cauchy–Dirichlet problem in bounded domains. Among those, the fractional porous medium equation ($10<\theta<$ 1, $21<q<$ 2) is studied in [19, 21, 22, 57], where well-posedness is proved in the class of (possibly sign-changing) “$H^{*}$-solutions”, which is an analogue of the $H^{-1}$ solution for $\theta=1$ to the fractional setting, by a direct application of the Brézis–Kōmura theory (see [23, 24]) and asymptotic behaviors of nonnegative $H^{*}$ solutions are also studied (see also [16, 17]). Thanks to a Bénilan–Crandall type estimate, the convergence of nonnegative $H^{*}$ solutions (after a rescaling) to a single asymptotic profile is proved. The study on the fractional fast diffusion equation ($10<\theta<$ 1, $q>2$) is still ongoing. We refer the reader to a pioneer work [49], where Hölder regularity and asymptotic behavior of nonnegative solutions are studied, as well as to a recent comprehensive study [18], where existence and uniqueness for nonnegative “weak dual solutions” are proved by means of the Brézis–Kōmura theory (to construct $H^{*}$ solutions first as an approximation) along with a Steklov averaging and a monotonicity argument (available only for nonnegative data), and where $L^{p}$-$L^{q}$ smoothing estimates and (finite-time) extinction estimates (in various norms) are further discussed for nonnegative weak dual solutions. Moreover, some of these works (see, e.g., [19, 18]) also cover more general nonlocal elliptic operators (e.g., spectral fractional Laplacian). On the other hand, concerning $10<\theta<$ 1, the study of (possibly) sign-changing weak solutions has not yet been fully pursued, and furthermore, full convergence to asymptotic profiles has never been proved even for nonnegative weak solutions to the fractional fast diffusion equation (cf. see [39] for the classical one $\theta=1$).
@@ 2.1. Fractional Sobolev spaces
$\mathcal{X}_{\theta}(\Omega):=\left\{w\in H^{\theta}(\mathbb{R}^{d})\colon w=0\ \text{ a.e.~{}on }\mathbb{R}^{d}\setminus\Omega\right\}\quad\text{ for }\ \theta\in(0,1).$
In what follows, we shall use the same letter $f$ to denote the restriction of $f\in\mathcal{X}_{\theta}(\Omega)$ to $\Omega$ when no confusion can arise. Moreover, we may also use the same notation for the zero extension to $\mathbb{R}^{d}$ of $f:\Omega\to\mathbb{R}$.
A Poincaré type inequality is also available for the class $\mathcal{X}_{\theta}(\Omega)$, $10<\theta<$ 1 (see, e.g., [6][7, Proposition A.1]).
@@
We first recall fractional Sobolev embeddings (see, e.g., [36, Theorems 6.7, 6.9 and 7.2]).
@@
with compact and continuous canonical injections. Indeed, the compact embedding $L^{2}(\Omega)^{*}\hookrightarrow\mathcal{X}_{\theta}(\Omega)^{*}$ follows from the compact embedding $\mathcal{X}_{\theta}(\Omega)\hookrightarrow L^{2}(\Omega)$ along with Schauder’s theorem (see, e.g., [25, Theorem 6.4]).
@@ 3. Main results
Corollary 3.7 (Decay and extinction estimates for energy solutions (cf. [18, §2.6][49])).
@@
Theorem 3.8 (Quasi-convergence to asymptotic profiles (cf. [49, Theorem 5.4])).
@@
The porous medium case (i.e., $21<q<$ 2 and $\phi\geq 0$) has already been discussed in [19, Theorem 2.3], and hence, the novelty of Theorem 3.9 may reside in the fast diffusion case (i.e., $2<q<2^{*}_{\theta}$ and either $\phi\geq 0$ or $q$ is even).
@@
In this section, we give a proof for Theorem 3.5. In §4.1, we shall briefly prove the uniqueness part (i) in a standard way. The main ingredients of the proof will be exhibited in §4.2, §4.3 and §4.4, where the additional regularity of the energy solution as well as the energy identity and inequalities will directly be proved within the weak formulation as in Definition 3.1 without relying on any classical parabolic theories (see, e.g., [51, 40, 52]) even for $\theta=1$; actually, they are no longer useful for the fractional case $10<\theta<$ 1. The $L^{1}$ contraction estimate and the Bénilan–Crandall estimate may be more or less standard, but they will finally be proved in §4.5 for the completeness.
@@ 4.2. Existence of energy solutions
for $w\in X$. Combining the coercivity with the convexity and lower semicontinuity of $I_{n}$ on $X$, we infer that $I_{n}$ admits a unique minimizer $u_{n+1}$ over $X$ (see, e.g., [25, Corollary 3.23]). Furthermore, the minimizer $u_{n+1}\in X$ solves the Euler-Lagrange equation,
@@
Finally, we derive additional regularity of the weak solution as well as energy inequalities. Since $W^{1,2}(0,T;L^{2}(\Omega))\subset C([0,T];L^{2}(\Omega))$, we also see that $w=|u|^{(q-2)/2}u\in C([0,T];L^{2}(\Omega))$. Moreover, we can deduce from Lemma A.1 that $u\in C([0,T];L^{q}(\Omega))$ and $\rho=|u|^{q-2}u\in C([0,T];L^{q^{\prime}}(\Omega))$, which along with $u\in L^{\infty}(0,T;\mathcal{X}_{\theta}(\Omega))$ and $|u|^{q-2}u\in L^{\infty}(0,T;L^{2}(\Omega))$ yield $u\in C_{\rm weak}([0,T];\mathcal{X}_{\theta}(\Omega))$ and $|u|^{q-2}u\in C_{\rm weak}([0,T];L^{2}(\Omega))$, respectively (see [53, Chap. 3, Lemma 8.1]). Furthermore, recalling the uniform estimates from (4.16)–(4.18) along with (4.27) and (4.28), for a.e. $t\in(0,T)$, we can derive, up to a (not relabeled) subsequence, that
@@
which along with $\|\partial_{t}(|u|^{(q-2)/2}u)(\cdot)\|_{L^{2}(\Omega)}\in L^{2}(0,\infty)$ and [27, Theorem 1.4.40] yields the desired conclusion.
@@ 4.3. Energy identity
Then it is well known that $\psi^{*}$ is also proper lower-semicontinuous and convex in $E^{*}$ (see, e.g., [25, §1.4]). Let $\partial_{E}\psi:E\to 2^{E^{*}}$ and $\partial_{E^{*}}\psi^{*}:E^{*}\to 2^{E}$ denote subdifferential operator of $\psi$ and $\psi^{*}$, respectively, that is,
@@
The lemma above is a variant of the well-known chain-rule formula for subdifferentials in convex analysis (cf. see [23, Lemma 3.3]). Indeed, one can prove it by combining the chain-rule formula [30, Lemma 4.1] (see the appendix of [3] for a proof) with the facts that $E^{**}$ is identified with $E$ by reflexivity and (4.32) is equivalent to $u(t)\in\partial_{E^{*}}\psi(\xi(t))$ for a.e. $t\in(a,b)$. ∎
@@ 4.4. For general initial data
Concerning $\theta=1$, these facts can be obtained by combining the conclusion in the $H^{-1}$ framework and the uniqueness. An analogous argument are also available for $10<\theta<$ 1 (cf. see [18, Theorem 2.3]).
@@ 4.5. $L^{1}$ contraction and Bénilan–Crandall estimate
This subsection is devoted to providing a proof for (iii) of Theorem 3.5, which follows the same line as [63] for $\theta=1$ and [49, 18] for $10<\theta<$ 1 but will also be exhibited for the completeness.
@@
for a.e. $t>0$, up to a subsequence, as $h\to 0$ (see, e.g., [27, Corollary 1.4.39]). Hence we infer from (4.44) that
@@
(see, e.g., [23, Appendix]). Hence recalling $w=|u|^{(q-2)/2}u\in W^{1,2}(0,\infty;L^{2}(\Omega))\subset W^{1,1}_{\rm loc}((0,\infty);L^{1}(\Omega))$ and employing Proposition 4.7, one can conclude that
@@ 5. Monotonicity of Rayleigh quotient and decay estimates
Let $0<T<t_{*}$ be fixed. The function $t\mapsto\|u(t)\|_{L^{q}(\Omega)}^{q}$ is absolutely continuous and uniformly away from 00 on $[0,T]$ (indeed, it is nonincreasing and $u(T)\not\equiv 0$). Moreover, due to (3.6) and (3.7) with $s=0$, the derivative $(\text{\rm d}/\text{\rm d}t)\|u(t)\|_{L^{q}(\Omega)}^{q}$ is uniformly bounded a.e. in $(0,T)$, and hence, the function $g(t):=1/\|u(t)\|_{L^{q}(\Omega)}^{2}$ is Lipschitz continuous on $[0,T]$. Thanks to (ii) of Theorem 3.5, the function $f(t):=\|u(t)\|_{\mathcal{X}_{\theta}(\Omega)}^{2}$ is nonincreasing on $[0,T]$ (and hence, of bounded variation in $(0,T)$) and right-continuous on $[0,T]$. Therefore ${\mathsf{R}}=fg$ is also right-continuous on $[0,T]$ and of bounded variation in $(0,T)$ (see [8, Proposition 3.2]). Therefore $\mathsf{R}$ is differentiable a.e. in $(0,T)$.
@@
We next prove the monotonicity of the Rayleigh quotient $t\mapsto\mathsf{R}(t)$ on $[0,\infty)$, which is still unclear, since $\mathsf{R}$ is of bounded variation only and may not be absolutely continuous. Since $f(t)=\|u(t)\|_{\mathcal{X}_{\theta}(\Omega)}^{2}$ is of bounded variation in $(0,T)$ and right-continuous on $[0,T]$, thanks to [8, Theorem 3.28] and Lebesgue’s decomposition theorem, it follows that, for $0<s<t<T$,
@@
where $\mathrm{D}f$ and $(\mathrm{D}f)_{s}$ denote the distributional derivative of $f$ represented as a finite Radon measure in $(0,T)$ and its singular part (with respect to the one-dimensional Lebesgue measure $\mathcal{L}^{1}$), respectively. Since $f$ is nonincreasing (see, e.g., [8, Corollary 3.29]), we see that
@@
which implies $(\mathrm{D}f)_{s}((s,t])\leq 0$ for any $(s,t]\subset(0,T)$. Since $f$ is of bounded variation and $g$ is Lipschitz continuous in $(0,T)$, by virtue of [8, Proposition 3.2] and Lebesgue’s decomposition theorem, we observe that
@@
where $\mathcal{L}^{1}$ denotes the one-dimensional Lebesgue measure. Therefore as ${\mathsf{R}}$ is of bounded variation in $(0,T)$ and right-continuous on $[0,T]$, we infer from [8, Theorem 3.28] and (5.3) that
@@ 6.1. Rescaled energy solutions
where $\mathrm{D}\phi$ and $(\mathrm{D}\psi)_{s}$ denote the distributional derivative of $\phi$ represented as a finite Radon measure in $(0,T)$ and the singular part (i.e., $(\mathrm{D}\psi)_{s}\perp\mathcal{L}^{1}$) of the distributional derivative $\mathrm{D}\psi$ of $\psi$, respectively, and $\mathcal{L}^{1}$ denotes the one-dimensional Lebesgue measure. As in the proof of Theorem 3.6, since $\psi$ is the product of a smooth positive function and a nonincreasing right-continuous function, we can also show that $(\mathrm{D}\psi)_{s}((s_{1},s_{2}])\leq 0$ for any $(s_{1},s_{2}]\subset(0,T)$. As $\phi=J(v(\cdot))$ is right-continuous on $[0,T]$ and of bounded variation in $(0,T)$, we infer from (6.2), (6.5) and [8, Theorem 3.28] that

### v2

@@ 1. Introduction
Most of studies on fractional nonlinear diffusion equations have been performed for the Cauchy problem, where the equations are posed in $\mathbb{R}^{d}$ (see, e.g., [37, 38, 69, 21, 35, 36, 15] and references therein). On the other hand, there have also been a limited number of contributions to the Cauchy–Dirichlet problem in bounded domains. Among those, the fractional porous medium equation ($10<\theta<$ 1, $21<q<$ 2) is studied in [20, 22, 23, 62], where well-posedness is proved in the class of (possibly sign-changing) “$H^{*}$-solutions”, which is an analogue of the $H^{-1}$ solution for $\theta=1$ to the fractional setting, by a direct application of the Brézis–Kōmura theory (see [26, 27]) and asymptotic behaviors of nonnegative $H^{*}$ solutions are also studied (see also [17, 18]). Thanks to a Bénilan–Crandall type estimate, the convergence of nonnegative $H^{*}$ solutions (after a rescaling) to a single asymptotic profile is proved. Moreover, convergence of (possibly) sign-changing weak solutions (after the rescaling) to sign-definite asymptotic profiles in the uniform norm are also proved under certain variational criteria for initial energy in [44], whose strategy is based on the paper [25] concerned with the (non-fractional) porous medium equation (see also [6, Remark 3.3] for some instability results). The study on the fractional fast diffusion equation ($10<\theta<$ 1, $q>2$) is still ongoing. We refer the reader to a pioneer work [54], where Hölder regularity and asymptotic behavior of nonnegative solutions are studied, as well as to a recent comprehensive study [19], where existence and uniqueness for nonnegative “weak dual solutions” are proved by means of the Brézis–Kōmura theory (to construct $H^{*}$ solutions first as an approximation) along with a Steklov averaging and a monotonicity argument (available only for nonnegative data), and where $L^{p}$-$L^{q}$ smoothing estimates and (finite-time) extinction estimates (in various norms) are further discussed for nonnegative weak dual solutions. Moreover, some of these works (see, e.g., [20, 19]) also cover more general nonlocal elliptic operators (e.g., spectral fractional Laplacian). On the other hand, concerning $10<\theta<$ 1, the study of (possibly) sign-changing weak solutions has not yet been fully pursued, and furthermore, full convergence to asymptotic profiles has never been proved even for nonnegative weak solutions to the fractional fast diffusion equation (cf. see [43] for the classical one $\theta=1$).
@@ 2.1. Fractional Sobolev spaces
$\mathcal{X}_{\theta}(\Omega):=\left\{w\in H^{\theta}(\mathbb{R}^{d})\colon w=0\ \text{ a.e.~{}on }\mathbb{R}^{d}\setminus\Omega\right\}\quad\text{ for }\ \theta\in(0,1),$
for which we also refer the reader to [24, 41] and references therein. In what follows, we shall use the same letter $f$ to denote the restriction of $f\in\mathcal{X}_{\theta}(\Omega)$ to $\Omega$ when no confusion can arise. Moreover, we may also use the same notation for the zero extension to $\mathbb{R}^{d}$ of $f:\Omega\to\mathbb{R}$.
A Poincaré type inequality is also available for the class $\mathcal{X}_{\theta}(\Omega)$, $10<\theta<$ 1 (see, e.g., [7][8, Proposition A.1]).
@@
We first recall fractional Sobolev embeddings (see, e.g., [39, Theorems 6.7, 6.9 and 7.2]).
@@
with compact and continuous canonical injections. Indeed, the compact embedding $L^{2}(\Omega)^{*}\hookrightarrow\mathcal{X}_{\theta}(\Omega)^{*}$ follows from the compact embedding $\mathcal{X}_{\theta}(\Omega)\hookrightarrow L^{2}(\Omega)$ along with Schauder’s theorem (see, e.g., [28, Theorem 6.4]).
@@ 3. Main results
Corollary 3.7 (Decay and extinction estimates for energy solutions (cf. [19, §2.6][54])).
@@
Theorem 3.8 (Quasi-convergence to asymptotic profiles (cf. [54, Theorem 5.4])).
@@
The porous medium case (i.e., $21<q<$ 2 and $\phi\geq 0$) has already been discussed in [20, Theorem 2.3] and [44, Theorem 1.1] for sign-definite and (possibly) sign-changing rescaled weak solutions $v=v(x,s)$, respectively, and hence, the novelty of Theorem 3.9 may reside in the fast diffusion case (i.e., $2<q<2^{*}_{\theta}$ and either $\phi\geq 0$ or $q$ is even).
@@
In this section, we give a proof for Theorem 3.5. In §4.1, we shall briefly prove the uniqueness part (i) in a standard way. The main ingredients of the proof will be exhibited in §4.2, §4.3 and §4.4, where the additional regularity of the energy solution as well as the energy identity and inequalities will directly be proved within the weak formulation as in Definition 3.1 without relying on any classical parabolic theories (see, e.g., [56, 45, 57]) even for $\theta=1$; actually, they are no longer useful for the fractional case $10<\theta<$ 1. The $L^{1}$ contraction estimate and the Bénilan–Crandall estimate may be more or less standard (see, e.g., [68, 38]), but they will finally be proved in §4.5 for the completeness.
@@ 4.2. Existence of energy solutions
for $w\in X$. Combining the coercivity with the convexity and lower semicontinuity of $I_{n}$ on $X$, we infer that $I_{n}$ admits a unique minimizer $u_{n+1}$ over $X$ (see, e.g., [28, Corollary 3.23]). Furthermore, the minimizer $u_{n+1}\in X$ solves the Euler-Lagrange equation,
@@
Finally, we derive additional regularity of the weak solution as well as energy inequalities. Since $W^{1,2}(0,T;L^{2}(\Omega))\subset C([0,T];L^{2}(\Omega))$, we also see that $w=|u|^{(q-2)/2}u\in C([0,T];L^{2}(\Omega))$. Moreover, we can deduce from Lemma A.1 that $u\in C([0,T];L^{q}(\Omega))$ and $\rho=|u|^{q-2}u\in C([0,T];L^{q^{\prime}}(\Omega))$, which along with $u\in L^{\infty}(0,T;\mathcal{X}_{\theta}(\Omega))$ and $|u|^{q-2}u\in L^{\infty}(0,T;L^{2}(\Omega))$ yield $u\in C_{\rm weak}([0,T];\mathcal{X}_{\theta}(\Omega))$ and $|u|^{q-2}u\in C_{\rm weak}([0,T];L^{2}(\Omega))$, respectively (see [58, Chap. 3, Lemma 8.1]). Furthermore, recalling the uniform estimates from (4.16)–(4.18) along with (4.27) and (4.28), for a.e. $t\in(0,T)$, we can derive, up to a (not relabeled) subsequence, that
@@
which along with $\|\partial_{t}(|u|^{(q-2)/2}u)(\cdot)\|_{L^{2}(\Omega)}\in L^{2}(0,\infty)$ and [30, Theorem 1.4.40] yields the desired conclusion.
@@ 4.3. Energy identity
Then it is well known that $\psi^{*}$ is also proper lower-semicontinuous and convex in $E^{*}$ (see, e.g., [28, §1.4]). Let $\partial_{E}\psi:E\to 2^{E^{*}}$ and $\partial_{E^{*}}\psi^{*}:E^{*}\to 2^{E}$ denote subdifferential operator of $\psi$ and $\psi^{*}$, respectively, that is,
@@
The lemma above is a variant of the well-known chain-rule formula for subdifferentials in convex analysis (cf. see [26, Lemma 3.3]). Indeed, one can prove it by combining the chain-rule formula [33, Lemma 4.1] (see the appendix of [3] for a proof) with the facts that $E^{**}$ is identified with $E$ by reflexivity and (4.32) is equivalent to $u(t)\in\partial_{E^{*}}\psi(\xi(t))$ for a.e. $t\in(a,b)$. ∎
@@ 4.4. For general initial data
Concerning $\theta=1$, these facts can be obtained by combining the conclusion in the $H^{-1}$ framework and the uniqueness. An analogous argument are also available for $10<\theta<$ 1 (cf. see [19, Theorem 2.3]).
@@ 4.5. $L^{1}$ contraction and Bénilan–Crandall estimate
This subsection is devoted to providing a proof for (iii) of Theorem 3.5, which follows the same line as [68] for $\theta=1$ and [38, 54, 19] for $10<\theta<$ 1 but will also be exhibited for the completeness.
@@
for a.e. $t>0$, up to a subsequence, as $h\to 0$ (see, e.g., [30, Corollary 1.4.39]). Hence we infer from (4.44) that
@@
(see, e.g., [26, Appendix]). Hence recalling $w=|u|^{(q-2)/2}u\in W^{1,2}(0,\infty;L^{2}(\Omega))\subset W^{1,1}_{\rm loc}((0,\infty);L^{1}(\Omega))$ and employing Proposition 4.7, one can conclude that
@@ 5. Monotonicity of Rayleigh quotient and decay estimates
Let $0<T<t_{*}$ be fixed. The function $t\mapsto\|u(t)\|_{L^{q}(\Omega)}^{q}$ is absolutely continuous and uniformly away from 00 on $[0,T]$ (indeed, it is nonincreasing and $u(T)\not\equiv 0$). Moreover, due to (3.6) and (3.7) with $s=0$, the derivative $(\text{\rm d}/\text{\rm d}t)\|u(t)\|_{L^{q}(\Omega)}^{q}$ is uniformly bounded a.e. in $(0,T)$, and hence, the function $g(t):=1/\|u(t)\|_{L^{q}(\Omega)}^{2}$ is Lipschitz continuous on $[0,T]$. Thanks to (ii) of Theorem 3.5, the function $f(t):=\|u(t)\|_{\mathcal{X}_{\theta}(\Omega)}^{2}$ is nonincreasing on $[0,T]$ (and hence, of bounded variation in $(0,T)$) and right-continuous on $[0,T]$. Therefore ${\mathsf{R}}=fg$ is also right-continuous on $[0,T]$ and of bounded variation in $(0,T)$ (see [9, Proposition 3.2]). Therefore $\mathsf{R}$ is differentiable a.e. in $(0,T)$.
@@
We next prove the monotonicity of the Rayleigh quotient $t\mapsto\mathsf{R}(t)$ on $[0,\infty)$, which is still unclear, since $\mathsf{R}$ is of bounded variation only and may not be absolutely continuous. Since $f(t)=\|u(t)\|_{\mathcal{X}_{\theta}(\Omega)}^{2}$ is of bounded variation in $(0,T)$ and right-continuous on $[0,T]$, thanks to [9, Theorem 3.28] and Lebesgue’s decomposition theorem, it follows that, for $0<s<t<T$,
@@
where $\mathrm{D}f$ and $(\mathrm{D}f)_{s}$ denote the distributional derivative of $f$ represented as a finite Radon measure in $(0,T)$ and its singular part (with respect to the one-dimensional Lebesgue measure $\mathcal{L}^{1}$), respectively. Since $f$ is nonincreasing (see, e.g., [9, Corollary 3.29]), we see that
@@
which implies $(\mathrm{D}f)_{s}((s,t])\leq 0$ for any $(s,t]\subset(0,T)$. Since $f$ is of bounded variation and $g$ is Lipschitz continuous in $(0,T)$, by virtue of [9, Proposition 3.2] and Lebesgue’s decomposition theorem, we observe that
@@
where $\mathcal{L}^{1}$ denotes the one-dimensional Lebesgue measure. Therefore as ${\mathsf{R}}$ is of bounded variation in $(0,T)$ and right-continuous on $[0,T]$, we infer from [9, Theorem 3.28] and (5.3) that
@@ 6.1. Rescaled energy solutions
where $\mathrm{D}\phi$ and $(\mathrm{D}\psi)_{s}$ denote the distributional derivative of $\phi$ represented as a finite Radon measure in $(0,T)$ and the singular part (i.e., $(\mathrm{D}\psi)_{s}\perp\mathcal{L}^{1}$) of the distributional derivative $\mathrm{D}\psi$ of $\psi$, respectively, and $\mathcal{L}^{1}$ denotes the one-dimensional Lebesgue measure. As in the proof of Theorem 3.6, since $\psi$ is the product of a smooth positive function and a nonincreasing right-continuous function, we can also show that $(\mathrm{D}\psi)_{s}((s_{1},s_{2}])\leq 0$ for any $(s_{1},s_{2}]\subset(0,T)$. As $\phi=J(v(\cdot))$ is right-continuous on $[0,T]$ and of bounded variation in $(0,T)$, we infer from (6.2), (6.5) and [9, Theorem 3.28] that